# The consistency strength of determinacy when all sets are universally Baire

### v2

@@ 5.6. Characterization via operators
For the other direction, suppose $E$ is the extender derived from $k\colon S_{i}(\mathcal{M}|\kappa_{i})\rightarrow S_{i}(\mathcal{M}|\xi(E))$. Let $\mathcal{Q}$ be a $\kappa_{i}$-good universal mouse over $\mathcal{M}|\kappa_{i}$ and let $\mathcal{Q}^{*}$ be a $\kappa_{i}$-good universal mouse over $\mathcal{M}|\xi(E)$. Consider the corresponding direct limits $\mathcal{R}^{\mathcal{Q}}$ and $\mathcal{R}^{\mathcal{Q}^{*}}$ of suitable premice. Both, $\mathcal{R}^{\mathcal{Q}}$ and $\mathcal{R}^{\mathcal{Q}^{*}}$, are $\Sigma_{i}$-iterates of $\mathcal{P}^{i}$. Moreover, $\mathcal{R}^{\mathcal{Q}}$ is a point in the direct limit system giving rise to $\mathcal{R}^{\mathcal{Q}^{*}}$. In fact, $\mathcal{R}^{\mathcal{Q}^{*}}$ is a $(\Sigma_{i})_{\mathcal{R}^{\mathcal{Q}}}$-iterate of $\mathcal{R}^{\mathcal{Q}}$. Note that $\mathcal{M}|\delta_{i+1}$ is a $\kappa_{i}$-good universal mouse over $\mathcal{M}|\kappa_{i}$ with $\mathcal{R}^{\mathcal{M}|\delta_{i+1}}=\mathcal{M}_{\infty}^{i}$. So we can pick $\mathcal{Q}=\mathcal{M}|\delta_{i+1}$ and the iteration from $\mathcal{R}^{\mathcal{Q}}=\mathcal{M}_{\infty}^{i}$ to $\mathcal{R}^{\mathcal{Q}^{*}}$ is according to $(\Sigma_{i})_{\mathcal{R}^{\mathcal{Q}}}$. Therefore, we obtain $\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i,-}=\pi^{\Sigma^{i,\infty}}_{\mathcal{M}_{\infty}^{i,-},\pi_{E}(\mathcal{M}_{\infty}^{i,-})}$, as desired. ∎
An important advantage of the characterization of extenders via operators is that it can be used to prove that the extenders that can be characterized this way are easily definable. This is made precise in the following lemma (cf. [13, Lemma 2.24 and Lemma 2.25]). The proof is analogous to the proofs in [13], so we leave it to the reader.
@@
Let $\xi$ be a cardinal in $\mathcal{M}$ with $(\kappa_{i}^{+})^{\mathcal{M}}<\xi<\delta_{i+1}$ for some $i<\omega$. Then
$S_{i}(\mathcal{M}|\xi)\in M_{1}^{\#,\Sigma_{i}}(\mathcal{M}|\xi).$
Moreover, if $E$ is an $\mathcal{M}|\delta_{i+1}$-extender with critical point $\kappa_{i}$ that is characterized via the operator $S_{i}$, then
$E\in M_{1}^{\#,\Sigma_{i}}(\operatorname{Ult}(\mathcal{M},E)|\zeta),$
where $\zeta$ is the least $i$-weak cutpoint above $\kappa_{i}$ in $\operatorname{Ult}(\mathcal{M},E)$.
@@ 6.1. Iterability
Our concept of an almost linear iteration tree is a generalization of the one in [20] as we distinguish two different types of extenders – with and without a background – and allow arbitrary iteration trees for backgrounded extenders. We make this more precise now and use the notation from [26] for iteration trees $\mathcal{T}$ and the associated tree order $T$.
@@
The following lemma uses similar ideas as the proof of [20, Lemma 2.4].
@@
Let $\mathcal{N}$ be the result of an $\mathcal{M}$-construction in some translatable structure. Then any normal iteration tree $\mathcal{T}$ on $\mathcal{N}$ is almost linear.
@@
Analogous to the proof of [20, Lemma 2.4], we start with the following claim.
@@
Suppose the claim does not hold and let $F$ be a strongness witness on the sequence of $\mathcal{M}_{\lambda}^{\mathcal{U}}$ with $\kappa_{E}\leq\kappa_{F}<\lambda_{E}\leq\lambda_{F}$. Then it is not hard to see that $\rho_{\omega}(J_{\lambda_{F}}^{\mathcal{M}_{\lambda}^{\mathcal{U}}})\geq\lambda_{E}$, cf. the subclaim in the proof of [20, Lemma 2.4, Claim 1]. This implies that $F|\xi\in J_{\lambda_{F}}^{\mathcal{M}_{\lambda}^{\mathcal{U}}}$ for all $\kappa_{F}<\xi<\lambda_{E}$. In particular, $\kappa_{F}$ is $\lambda_{E}$-strong in $\mathcal{M}_{\lambda}^{\mathcal{U}}$. Therefore, $\kappa_{F}$ is $\lambda_{E}$-strong in $\mathcal{M}_{i+1}^{\mathcal{U}}$. This yields a contradiction as $\mathcal{M}_{i+1}^{\mathcal{U}}$ is an ultrapower by $E$, so there is no extender with critical point between $\kappa_{E}$ and $\lambda_{E}$ that overlaps $\lambda_{E}$ in $\mathcal{M}_{i+1}^{\mathcal{U}}$. ∎
@@
Using this observation, a short argument as in the proof of [20, Lemma 2.4] yields that any normal iteration tree $\mathcal{T}$ on $\mathcal{N}$ is almost linear. ∎
@@
Let $i<\omega$ and let $\bar{\mathcal{W}}$ be the transitive collapse of a hull of $\mathcal{W}|\Omega$ for some sufficiently large $\Omega$ such that $\mathcal{W}|\kappa_{i}\subseteq\bar{\mathcal{W}}$. Let
$\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}$
be the uncollapse embedding, $\bar{\mathcal{M}}$ the collapse of $\mathcal{M}|\Omega$, and $(\bar{\kappa}_{j},\bar{\delta}_{j}\mid j<\omega)$ the preimages of $(\kappa_{j},\delta_{j}\mid j<\omega)$ (so, in particular, $\bar{\kappa}_{j}=\kappa_{j}$ and $\bar{\delta}_{j}=\delta_{j}$ for $j\leq i$). Then $\bar{\mathcal{M}}$ is $(\delta_{i+1},\delta_{i+1})$-iterable with respect to extenders with critical point above $\kappa_{i}$.
@@
Let $i<\omega$, $\bar{\mathcal{W}}$ and $\bar{\mathcal{M}}$ be as in the statement of the lemma. For notational simplicity we write $\mathcal{M}$ and $\mathcal{W}$ instead of $\mathcal{M}|\Omega$ and $\mathcal{W}|\Omega$ as it will be clear from the context if we refer to the proper class model or the restriction to $\Omega$. We recursively describe an iteration strategy $\Lambda$ for $\bar{\mathcal{M}}$ acting on iteration trees $\mathcal{T}$ of length ${<}\delta_{i+1}$ which use extenders with critical point above $\kappa_{i}$. The recursive construction will depend on what type of extenders get used – backgrounded extenders or non-backgrounded strongness witnesses. For notational simplicity, we describe the iteration strategy $\Lambda$ only for normal trees $\mathcal{T}$ and leave it to the reader to generalize the construction to stacks of normal trees. Simultaneously to the construction of $\Lambda$, we define a realization of $\mathcal{T}$ into a backgrounded iteration of $\mathcal{M}$. That means, for every normal iteration tree $\mathcal{T}$ on $\bar{\mathcal{M}}$ we recursively define a stack of iteration trees $\vec{\mathcal{U}}=(\mathcal{U}_{\alpha}\mid\alpha<\gamma)$ on $\mathcal{M}$ and an embedding
@@
Now we turn to the recursive definition of $\Lambda$ together with realizations into a backgrounded iteration of $\mathcal{M}$. Let $\mathcal{T}$ be a normal iteration tree of limit length $\lambda<\delta_{i+1}$ on $\bar{\mathcal{M}}$ according to $\Lambda$ and recall that $\mathcal{T}$ is almost linear by Lemma 6.2. We distinguish the following cases.
@@
Finally, we define $\Lambda(\mathcal{T})$ to be the canonical branch through $\mathcal{T}$ (or more precisely through $\mathcal{T}_{2}$) obtained from the iteration strategy for the background model $\mathcal{W}$ by lifting the stack of the two trees $(\iota_{0}\mathcal{T}_{1}\upharpoonright\mathcal{M}^{1},\iota_{\xi+1}\mathcal{T}_{2})$ to a stack of two trees on $\mathcal{W}$.
@@
By Lemma 6.2, $\mathcal{T}$ is an almost linear iteration tree. Therefore, there is a canonical cofinal branch $b$ through $\mathcal{T}$ in this case and we can let $\Lambda(\mathcal{T})=b$. To continue, we can consider the tree $\mathcal{T}^{\frown}b$ of length $\lambda+1$ and let $\iota_{\lambda}$ be the limit of $\iota_{\xi}$ for $\xi\in b$.
@@ 6.2. Comparison against a hull
The following lemma will be crucial in the proof that the construction of $\mathcal{M}$ converges and all $\kappa_{i}$ for $i<\omega$ are fully strong in $\mathcal{M}$. We inductively assume that all $\kappa_{j}$ for $j<i$ are strong in $\mathcal{M}$ and use the lemma to prove that $\kappa_{i}$ is strong in $\mathcal{M}$ in Lemma 6.6.
@@
Let $i<\omega$ and suppose that all cardinals $\kappa_{j}$ for $j<i$ are fully strong in $\mathcal{M}$. Let $\bar{\mathcal{W}}$ be the transitive collapse of a hull of $\mathcal{W}|\Omega$ for some sufficiently large $\Omega$ such that $\mathcal{W}|\kappa_{i}\subseteq\bar{\mathcal{W}}$ and $|\bar{\mathcal{W}}|=\kappa_{i}$. Let
$\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}$
be the uncollapse embedding, $\bar{\mathcal{M}}$ the collapse of $\mathcal{M}|\Omega$, and $(\bar{\kappa}_{j},\bar{\delta}_{j}\mid j<\omega)$ the preimages of $(\kappa_{j},\delta_{j}\mid j<\omega)$ (so, in particular, $\bar{\kappa}_{j}=\kappa_{j}$ and $\bar{\delta}_{j}=\delta_{j}$ for $j\leq i$). Then there is an iterate $\mathcal{M}^{*}$ of $\bar{\mathcal{M}}$ such that $\bar{\mathcal{M}}$$\mathcal{M}^{*}$ is non-dropping, $\mathcal{M}^{*}=\mathcal{M}_{\chi}$ for some level $\mathcal{M}_{\chi}$ in the construction of $\mathcal{M}$, and if
@@
By Lemma 6.3 we have that $\bar{\mathcal{M}}$ is $(\delta_{i+1},\delta_{i+1})$-iterable with respect to extenders with critical point above $\kappa_{i}$ and we can compare it against the construction of $\mathcal{M}$ (cf., e.g. [17] for such a comparison argument for $K^{c}$ constructions). We have |¯$|\bar{\mathcal{M}}|<\delta_{i+1}$ and we compare $\bar{\mathcal{M}}$ against models $\mathcal{M}_{\xi}$ of the construction of $\mathcal{M}$ with $|\mathcal{M}_{\xi}|<\delta_{i+1}$. We will see below that the comparison against the construction of $\mathcal{M}$ will finish successfully before reaching models $\mathcal{M}_{\xi}$ of size larger than $\delta_{i+1}$. So the iteration trees we consider on $\bar{\mathcal{M}}$ will have length at most $\max(|\bar{\mathcal{M}}|,|\mathcal{M}_{\xi}|)^{+}<\delta_{i+1}$ for ordinals $\xi$ with $|\mathcal{M}_{\xi}|<\delta_{i+1}$. In addition, we will argue in Claim 2 below that all extenders used in this comparison on the $\bar{\mathcal{M}}$-side have critical point above $\kappa_{i}$, so the iterability given by Lemma 6.3 suffices for $\bar{\mathcal{M}}$. Moreover, we start with arguing that no disagreement is caused by the $\mathcal{M}$-side, so the construction of $\mathcal{M}$ does not move in the comparison process. This also shows that we do not need iterability for $\mathcal{M}$ itself for the comparison argument.
@@
Let $(\mathcal{M}_{\xi},\mathcal{N}_{\xi}\mid\xi\leq\Omega)$ be the models in the construction of $\mathcal{M}$ and suppose toward a contradiction that the construction of $\mathcal{M}$ does move in the comparison. Let $E$ be the first extender on some $\mathcal{M}_{\xi}$ in the construction of $\mathcal{M}$ that is used. Write $\mathcal{M}^{*}$ for the iterate of $\bar{\mathcal{M}}$ we have obtained up to this stage in the comparison and $\mathcal{T}$ for the corresponding iteration tree of length $\lambda+1$ for some ordinal $\lambda$, i.e., $\mathcal{M}_{\lambda}^{\mathcal{T}}=\mathcal{M}^{*}$. As the construction of $\mathcal{M}$ wins the comparison, there is no drop on the main branch of the $\bar{\mathcal{M}}$-side and we have an elementary embedding $i=i_{0\lambda}^{\mathcal{T}}\colon\bar{\mathcal{M}}\rightarrow\mathcal{M}^{*}$.
It is a standard argument that fully backgrounded extenders with critical point above $\kappa_{i}$ do not move in such a comparison (see for example [12, Lemma 2.11] or [17]), so we can suppose that $\operatorname{crit}(E)\in\{\kappa_{j}\mid j<\omega\}$. As |¯$|\bar{\mathcal{M}}|<\delta_{i+1}$ and $|\mathcal{M}_{0}|<\delta_{i+1}$ their comparison lasts at most up to $\max(|\bar{\mathcal{M}}|,|\mathcal{M}_{0}|)^{+}<\delta_{i+1}$. Repeating this implies $\operatorname{crit}(E)=\kappa_{j}$ for some $j\leq i$ as $\bar{\mathcal{M}}$ cannot iterate up to $\delta_{i+1}<\kappa_{i+1}$.
Extenders on both sides: We start by arguing that if there is an extender $F$ in $\mathcal{M}^{*}$ with the same index as $E$ then $F$ is certified via the same tail strategy of $\Sigma_{j}$ as $F$ and hence $E=F$, contradicting the assumption that $E$ is used because of a disagreement. To see that $F$ is certified by the same strategy certifying $E$, we consider different cases depending on how the iteration from $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ looks like. Let $\bar{\Sigma}$ denote the iteration strategy of $\bar{\mathcal{W}}$ obtained as the pullback of $\Sigma$.

### v3

@@ 5.6. Characterization via operators
For the other direction, suppose $E$ is the extender derived from $k\colon S_{i}(\mathcal{M}^{(i)}|\kappa_{i})\rightarrow S_{i}(\mathcal{M}^{(i)}|\xi(E))$. Let $\mathcal{Q}$ be a $\kappa_{i}$-good universal mouse over $\mathcal{M}^{(i)}|\kappa_{i}$ and let $\mathcal{Q}^{*}$ be a $\kappa_{i}$-good universal mouse over $\mathcal{M}^{(i)}|\xi(E)$. Consider the corresponding direct limits $\mathcal{R}^{\mathcal{Q}}$ and $\mathcal{R}^{\mathcal{Q}^{*}}$ of suitable premice. Both, $\mathcal{R}^{\mathcal{Q}}$ and $\mathcal{R}^{\mathcal{Q}^{*}}$, are $\Sigma_{i}$-iterates of $\mathcal{P}^{i}$. Moreover, $\mathcal{R}^{\mathcal{Q}}$ is a point in the direct limit system giving rise to $\mathcal{R}^{\mathcal{Q}^{*}}$. In fact, $\mathcal{R}^{\mathcal{Q}^{*}}$ is a $(\Sigma_{i})_{\mathcal{R}^{\mathcal{Q}}}$-iterate of $\mathcal{R}^{\mathcal{Q}}$. Note that $\mathcal{M}^{(i)}|\delta_{i+1}$ is a $\kappa_{i}$-good universal mouse over $\mathcal{M}^{(i)}|\kappa_{i}$ with $\mathcal{R}^{\mathcal{M}^{(i)}|\delta_{i+1}}=\mathcal{P}_{\infty}^{i}$. So we can pick $\mathcal{Q}=\mathcal{M}^{(i)}|\delta_{i+1}$ and the iteration from $\mathcal{R}^{\mathcal{Q}}=\mathcal{P}_{\infty}^{i}$ to $\mathcal{R}^{\mathcal{Q}^{*}}$ is according to $(\Sigma_{i})_{\mathcal{R}^{\mathcal{Q}}}$. Therefore, we obtain $\pi_{E}\upharpoonright\mathcal{P}_{\infty}^{i,-}=\pi^{\Sigma^{i,\infty}}_{\mathcal{P}_{\infty}^{i,-},\pi_{E}(\mathcal{P}_{\infty}^{i,-})}$, as desired. ∎
An important advantage of the characterization of extenders via operators is that it can be used to prove that the extenders that can be characterized this way are easily definable. This is made precise in the following lemma (cf. [14, Lemma 2.24 and Lemma 2.25]).
@@
Let $i<\omega$ and suppose $(\mathsf{IH})_{i}$. Let $\xi$ be a cardinal in $\mathcal{M}^{(i)}$ with $(\kappa_{i}^{+})^{\mathcal{M}^{(i)}}<\xi<\delta_{i+1}$. Then
$S_{i}(\mathcal{M}^{(i)}|\xi)\in M_{1}^{\#,\Sigma_{i}}(\mathcal{M}^{(i)}|\xi).$
Moreover, if $E$ is an $\mathcal{M}^{(i)}|\delta_{i+1}$-extender with critical point $\kappa_{i}$ that is characterized via the operator $S_{i}$, then
$E\in M_{1}^{\#,\Sigma_{i}}(\operatorname{Ult}(\mathcal{M}^{(i)},E)|\zeta),$
where $\zeta$ is the least $i$-weak cutpoint above $\kappa_{i}$ in $\operatorname{Ult}(\mathcal{M}^{(i)},E)$.
The proof of Lemma 5.23 is analogous to the proofs of Lemmas 2.24 and 2.25 in [14], so we leave it to the reader. The idea of proof is that, considering $\operatorname{Ult}(\mathcal{M}^{(i)},E)$ for $E$ an extender on the sequence of $\mathcal{M}^{(i)}$ such that $\xi$ is a cutpoint in $\operatorname{Ult}(\mathcal{M}^{(i)},E)$, we can make $\mathcal{M}^{(i)}|\xi$ generic for the extender algebra of $\mathcal{M}^{\prime}$ for some iteration $i\colon\mathcal{M}^{(i)}\rightarrow\mathcal{M}^{\prime}$. Using $\mathcal{P}$-constructions we can conclude that $S_{i}(\mathcal{M}^{(i)}|\xi)\in M_{1}^{\#,\Sigma_{i}}(\mathcal{M}^{(i)}|\xi).$ Note that $S_{i}(\mathcal{M}^{(i)}|\kappa_{i})=Hull_{1}^{S_{i}(\operatorname{Ult}(\mathcal{M}^{(i)},E)|\zeta)}(\mathcal{M}^{(i)}|\kappa_{i})$ and let $\sigma\colon S_{i}(\mathcal{M}^{(i)}|\kappa_{i})\rightarrow S_{i}(\operatorname{Ult}(\mathcal{M}^{(i)},E)|\zeta)$ be the uncollapse map. The argument for the first part of Lemma 5.23 sketched above also shows $\sigma\in M_{1}^{\#,\Sigma_{i}}(\operatorname{Ult}(\mathcal{M}^{(i)},E)|\zeta)$. This yields $E\in M_{1}^{\#,\Sigma_{i}}(\operatorname{Ult}(\mathcal{M}^{(i)},E)|\zeta)$ as $E$ can be derived from $\sigma$.
@@ 6.1. Iterability
Our concept of an almost linear iteration tree is a generalization of the one in [21] as we distinguish two different types of extenders – with and without a background – and allow arbitrary iteration trees for backgrounded extenders. We make this more precise now and use the notation from [28] for iteration trees $\mathcal{T}$ and the associated tree order $T$. Lemma 6.2 will show that in our setting all normal iteration trees are almost linear on strongness witnesses.
@@
The following lemma uses similar ideas as the proof of [21, Lemma 2.4].
@@
Let $\mathcal{N}$ be the result of an $\mathcal{M}$-construction in some translatable structure. Then any normal iteration tree $\mathcal{T}$ on $\mathcal{N}$ is almost linear on strongness witnesses.
@@
Analogous to the proof of [21, Lemma 2.4], we start with the following claim.
@@
Suppose the claim does not hold and let $F$ be a strongness witness on the sequence of $\mathcal{M}_{\lambda}^{\mathcal{U}}$ with $\kappa_{E}\leq\kappa_{F}<\lambda_{E}\leq\lambda_{F}$. Then it is not hard to see that $\rho_{\omega}(J_{\lambda_{F}}^{\mathcal{M}_{\lambda}^{\mathcal{U}}})\geq\lambda_{E}$, cf. the subclaim in the proof of [21, Lemma 2.4, Claim 1]. This implies that $F|\xi\in J_{\lambda_{F}}^{\mathcal{M}_{\lambda}^{\mathcal{U}}}$ for all $\kappa_{F}<\xi<\lambda_{E}$. In particular, $\kappa_{F}$ is $\lambda_{E}$-strong in $\mathcal{M}_{\lambda}^{\mathcal{U}}$. Therefore, $\kappa_{F}$ is $\lambda_{E}$-strong in $\mathcal{M}_{i+1}^{\mathcal{U}}$. This yields a contradiction as $\mathcal{M}_{i+1}^{\mathcal{U}}$ is an ultrapower by $E$, so there is no extender with critical point between $\kappa_{E}$ and $\lambda_{E}$ that overlaps $\lambda_{E}$ in $\mathcal{M}_{i+1}^{\mathcal{U}}$. ∎
@@
Using this observation, a short argument as in the proof of [21, Lemma 2.4] yields that any normal iteration tree $\mathcal{T}$ on $\mathcal{N}$ is almost linear on strongness witnesses. ∎
The next lemma proves $(\mathsf{IH}.1)_{i+1}$ from $(\mathsf{IH})_{i}$.
@@
Let $i<\omega$, suppose $(\mathsf{IH})_{i}$ and let
$\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}|\Omega$
for some sufficiently large [^31] Note that we have not shown at this point that the construction of $\mathcal{M}^{(i+1)}$ converges. In case $\mathcal{M}^{(i+1)}\cap\operatorname{Ord}<\operatorname{Ord}$ pick $\Omega>\mathcal{M}^{(i+1)}\cap\operatorname{Ord}$ and let $\mathcal{M}^{(i+1)}|\Omega=\mathcal{M}^{(i+1)}$. $\Omega$ be such that letting $\operatorname{crit}(\pi)=\nu$, $\mathcal{W}|\kappa_{i}\subseteq\bar{\mathcal{W}}$, $(\kappa_{j},\delta_{j}\mid j<\omega)\in\operatorname{rng}(\pi)$, the sequence of models of the construction of $\mathcal{M}^{(i+1)}|\Omega$ is in the range of $\pi$, $\nu\in(\kappa_{i},\delta_{i+1})$ is an inaccessible cardinal in $\mathcal{W}$, and $\pi(\nu)=\delta_{i+1}$. Let $\bar{\mathcal{M}}$ be the collapse of $\mathcal{M}^{(i+1)}|\Omega$. Then $\bar{\mathcal{M}}$ is $(\delta_{i+1},\delta_{i+1})$-iterable with respect to extenders with critical point above $\kappa_{i}$.
@@
Let $i<\omega$ and $\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}|\Omega$ be as in the statement of the lemma. Let $(\bar{\kappa}_{j},\bar{\delta}_{j}\mid j<\omega)$ be the preimages of $(\kappa_{j},\delta_{j}\mid j<\omega)$ under $\pi$ (so, in particular, $\bar{\kappa}_{j}=\kappa_{j}$ and $\bar{\delta}_{j}=\delta_{j}$ for $j\leq i$). For notational simplicity we write $\mathcal{M}$ and $\mathcal{W}$ instead of $\mathcal{M}^{(i+1)}|\Omega$ and $\mathcal{W}|\Omega$ as it will be clear from the context if we refer to the proper class model or the restriction to $\Omega$ and suppressing the superscript ${(i+1)}$ simplifies the notation. We recursively describe an iteration strategy $\Lambda$ for $\bar{\mathcal{M}}$ acting on iteration trees $\mathcal{T}$ of length ${<}\delta_{i+1}$ which use extenders with critical point above $\kappa_{i}$. The recursive construction will depend on what type of extenders get used – backgrounded extenders or non-backgrounded strongness witnesses. For notational simplicity, we describe the iteration strategy $\Lambda$ only for normal trees $\mathcal{T}$ and leave it to the reader to generalize the construction to stacks of normal trees. Simultaneously to the construction of $\Lambda$, we define a realization of $\mathcal{T}$ into a backgrounded iteration of $\mathcal{M}$. That means, for every normal iteration tree $\mathcal{T}$ on $\bar{\mathcal{M}}$ we recursively define a stack of iteration trees $\vec{\mathcal{U}}=(\mathcal{U}_{\alpha}\mid\alpha<\gamma)$ on $\mathcal{M}$ and an embedding
@@
Now we turn to the recursive definition of $\Lambda$ together with realizations into a backgrounded iteration of $\mathcal{M}$. Let $\mathcal{T}$ be a normal iteration tree of limit length $\lambda<\delta_{i+1}$ on $\bar{\mathcal{M}}$ according to $\Lambda$ and recall that $\mathcal{T}$ is almost linear on strongness witnesses by Lemma 6.2. We distinguish the following cases.
@@
Finally, we define $\Lambda(\mathcal{T})$ to be the canonical branch through $\mathcal{T}$ (or more precisely through $\mathcal{T}_{2}$) obtained from the iteration strategy for the background model $\mathcal{W}$ by lifting the stack of the two trees $(\iota_{0}\mathcal{T}_{1}\upharpoonright\mathcal{M}^{1},\iota_{\xi+1}\mathcal{T}_{2})$ to a stack of two trees on $\mathcal{W}$. Note that $\Lambda(\mathcal{T})$ depends on our choice of realization map $\tau$ from Claim 1.
@@
By Lemma 6.2, $\mathcal{T}$ is almost linear on strongness witnesses. Therefore, there is a canonical cofinal branch $b$ through $\mathcal{T}$ in this case and we can let $\Lambda(\mathcal{T})=b$. To continue, we can consider the tree $\mathcal{T}^{\frown}b$ of length $\lambda+1$ and let $\iota_{\lambda}$ be the limit of $\iota_{\xi}$ for $\xi\in b$.
@@ 6.2. Comparison against a hull
The following lemma will be crucial in the proof that the construction of $\mathcal{M}$ converges and all $\kappa_{i}$ for $i<\omega$ are fully strong in $\mathcal{M}$.
@@
Let $i<\omega$, suppose $(\mathsf{IH})_{i}$ and let
$\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}|\Omega$
for some sufficiently large [^32] Note that we have not shown at this point that the construction of $\mathcal{M}^{(i+1)}$ converges. In case $\mathcal{M}^{(i+1)}\cap\operatorname{Ord}<\operatorname{Ord}$ pick $\Omega>\mathcal{M}^{(i+1)}\cap\operatorname{Ord}$ and let $\mathcal{M}^{(i+1)}|\Omega=\mathcal{M}^{(i+1)}$. $\Omega$ be such that letting $\operatorname{crit}(\pi)=\nu$, $\mathcal{W}|\kappa_{i}\subseteq\bar{\mathcal{W}}$, $(\kappa_{j},\delta_{j}\mid j<\omega)\in\operatorname{rng}(\pi)$, the sequence of models of the construction of $\mathcal{M}^{(i+1)}|\Omega$ is in the range of $\pi$, $\nu\in(\kappa_{i},\delta_{i+1})$ is an inaccessible cardinal in $\mathcal{W}$, and $\pi(\nu)=\delta_{i+1}$. Let $\bar{\mathcal{M}}$ be the collapse of $\mathcal{M}^{(i+1)}|\Omega$. Then there is an iterate $\mathcal{M}^{*}$ of $\bar{\mathcal{M}}$ such that $\bar{\mathcal{M}}$$\mathcal{M}^{*}$ is non-dropping, $\mathcal{M}^{*}=\mathcal{M}_{\chi}$ for some level $\mathcal{M}_{\chi}$ in the construction of $\mathcal{M}^{(i+1)}$, and if
@@
Let $(\bar{\kappa}_{j},\bar{\delta}_{j}\mid j<\omega)$ be the preimages of $(\kappa_{j},\delta_{j}\mid j<\omega)$ under $\pi$ (so, in particular, $\bar{\kappa}_{j}=\kappa_{j}$ and $\bar{\delta}_{j}=\delta_{j}$ for $j\leq i$). For notational simplicity we write $\mathcal{M}$ for $\mathcal{M}^{(i+1)}|\Omega$ and $\mathcal{W}$ for $\mathcal{W}|\Omega$. By Lemma 6.3 we have that $\bar{\mathcal{M}}$ is $(\delta_{i+1},\delta_{i+1})$-iterable with respect to extenders with critical point above $\kappa_{i}$ and we can compare its construction against the construction of $\mathcal{M}$ (cf., e.g. [18] or [29]). We have |¯$|\bar{\mathcal{M}}|<\delta_{i+1}$ and we compare the construction of $\bar{\mathcal{M}}$ against models $\mathcal{M}_{\xi}$ of the construction of $\mathcal{M}$ with $|\mathcal{M}_{\xi}|<\delta_{i+1}$. We will see below that the comparison against the construction of $\mathcal{M}$ will finish successfully before reaching models $\mathcal{M}_{\xi}$ of size larger than $\delta_{i+1}$. So the iteration trees we consider on models in the construction of $\bar{\mathcal{M}}$ will have length at most $\max(|\bar{\mathcal{M}}|,|\mathcal{M}_{\xi}|)^{+}<\delta_{i+1}$ for ordinals $\xi$ with $|\mathcal{M}_{\xi}|<\delta_{i+1}$. In addition, we will argue in Claim 2 below that all extenders used in this comparison on the $\bar{\mathcal{M}}$-side have critical point above $\kappa_{i}$, so the iterability given by Lemma 6.3 suffices for $\bar{\mathcal{M}}$. Moreover, we start with arguing that no disagreement is caused by the $\mathcal{M}$-side, so the construction of $\mathcal{M}$ does not move in the comparison process. This also shows that we do not need iterability for $\mathcal{M}$ itself for the comparison argument.
@@
Let $(\mathcal{M}_{\xi},\mathcal{N}_{\xi}\mid\xi\leq\Omega)$ be the models in the construction of $\mathcal{M}$ and suppose toward a contradiction that the construction of $\mathcal{M}$ does move in the comparison. Let $E$ be the first extender on some $\mathcal{M}_{\xi}$ in the construction of $\mathcal{M}$ that is used. Write $\mathcal{M}^{*}$ for the iterate of the $\bar{\mathcal{M}}$-side we have obtained up to this stage in the comparison and $\mathcal{T}$ for the corresponding iteration tree of length $\lambda+1$ for some ordinal $\lambda$, i.e., $\mathcal{M}_{\lambda}^{\mathcal{T}}=\mathcal{M}^{*}$. As the construction of $\mathcal{M}$ wins the comparison, there is no drop on the main branch of the $\bar{\mathcal{M}}$-side and we have an elementary embedding $i=i_{0\lambda}^{\mathcal{T}}\colon\bar{\mathcal{M}}_{\xi}\rightarrow\mathcal{M}^{*}$ for ¯$\bar{\mathcal{M}}_{\xi}$ a model in the construction of $\bar{\mathcal{M}}$.
It is a standard argument that fully backgrounded extenders with critical point above $\kappa_{i}$ do not move in such a comparison (see for example [13, Lemma 2.11] or [18]), so we can suppose that $\operatorname{crit}(E)\in\{\kappa_{j}\mid j<\omega\}$. $\bar{\mathcal{M}}$ cannot iterate up to $\delta_{i+1}<\kappa_{i+1}$ as, by the choice of $\bar{\mathcal{W}}$, $\bar{\delta}_{i+1}$ is not Woodin in $\mathcal{W}$ and (as we will argue below) the iteration of the construction of $\bar{\mathcal{M}}$ takes place above $\bar{\delta}_{i+1}$.
Extenders on both sides: We start by arguing that if there is an extender $F$ in $\mathcal{M}^{*}$ with the same index as $E$ then $F$ is certified via the same tail strategy of $\Sigma_{j}$ as $F$ and hence $E=F$, contradicting the assumption that $E$ is used because of a disagreement. To see that $F$ is certified by the same strategy certifying $E$, we consider different cases depending on how the iteration from the construction of $\bar{\mathcal{M}}$ to $\mathcal{M}^{*}$ looks like. Let $\bar{\Sigma}$ denote the iteration strategy of $\bar{\mathcal{W}}$ obtained as the pullback of $\Sigma$ and ¯$\bar{\mathcal{M}}_{\xi}$ the current model in the construction of $\bar{\mathcal{M}}$.